# Accounting for multiplicity in machine learning benchmark performance

### v4

@@ 4.2 Non-identical, independent classifiers
The Poisson Binomial distribution, see, e.g., Wang (1993), describes the sum of $n$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{n}$. The pmf for $K$ number of successes is
@@
where $G_{k}$ is all subsets that can be selected from {1, 2, …, n}.
In the multiple classifier context, we are interested in generalising Eq. 3.3 to the non-identical case.
$F(z)=P(C_{z}>0|m,n,\mathbf{\theta})=1-P(C_{z}=0|m,n,\mathbf{\theta}),$ (7)
where $P(C_{z}=0)$ corresponds to 00 classifiers with non-identical probabilities of success, $P(Z\leq z|\theta_{j})$. The cdf is
$F(z)$ $=1-\sum\limits_{{A\in G_{0}}}\prod\limits_{{i\in A}}P(Z\leq z|\theta_{i})\prod\limits_{{j\in A^{c}}}(1-P(Z\leq z|\theta_{j}))$
$=1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}.$ (8)
@@
$f(z)$ $=F(z)-F(z-1)$
$=1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}-\{1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z\text{-}1|\theta_{j})\big{)}\}$
$=\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z\text{-}1|\theta_{j})\big{)}-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}$ (9)
$=\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z\text{-1}}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}$
@@
We use a similar set-up as in Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, and probabilities of correct prediction $\theta=\{\theta_{1},\ldots,\theta_{m}\}$, where $\min\theta=0.875,\,\max\theta=0.90$, and they are equally spaced.
The simulated non-identical upper bound of the 95% confidence interval is $0.9177$, with $100,000$ repetitions. The $E(\hat{\theta}^{\,max}_{SOTA})$ is $0.9129$, with standard deviation of $0.0021$. The cdf and pmf are displayed in Fig. 5 and Fig. 6.
Cumulative distribution function for non-identical classifiers.
Figure 5:
Probability mass function for non-identical classifiers.
Figure 6:
@@ 4.3 Non-identical, dependent classifiers
The most complex, but also the most realistic case, is where classifiers are non-identical and dependent. An analytical solution is not known Hisakado et al. (2006), but simulations are easy. The results of Boland et al. (1989) are easily extended, and we have that
@@
Note that, since $P\leq 1$, it follows from 11 that
@@
and similarly for $0\leq P$ and 10.
We get the a the simulated dependent upper bound of the 0.95 confidence interval of $0.91733$, with $100,000$ repetitions.
@@ 5 Real world examples
As we have shown above with simulated data, the observed maximum is an overoptimistic estimate for SOTA. These results are based on known state-of-the-art, but faced with real world data, the problem to be solved is of opposite nature: The observed $\hat{\theta}^{\,max}$ is known, but the true SOTA must be estimated. In the public competition or public data set context there are many factors that come in to play, making the bias calculations everything but straightforward.
@@
In a public competition with $m$ participating teams, let $\hat{\theta}_{j}$, $j=1,\ldots,m$ be the estimated probability of correct prediction for each classifier based on the performance on an independent test set of size $n$. We will refer to these as the competition observations. $\hat{\theta}^{\,max}$ is the observed performance of the winning team. Among the different scenarios presented in Section 4, the non-identical dependent is the most realistic model. Whereas in Table 2 we simulated $E(\hat{\theta}^{\,max})$ from known $\theta_{1},\ldots,\theta_{m}$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\theta}_{1},\ldots,\hat{\theta}_{m}$. A ’natural’ approach is to find $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that gives $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$. That is, find a set of $\theta$’s so that the expected value of the observed maximum on a test set of size $n$ is equal to the competition observed maximum. We use a straightforward set-up, where the simulated $\hat{\tilde{\theta}}_{j}$ is the mean value of $n$ realisations of $Y_{j}$, where $P(Y_{j})=\tilde{\theta}_{j}$. The simulation is repeated $10,000$ times, and for each repetition the observed maximum value is recorded. $E(\hat{\tilde{\theta}}^{\,\max})$ is the mean value over the repetitions.
In the studies in Section 4, the $\theta_{1},\ldots,\theta_{m}$ were chosen to be equally spaced, but this does not align with the competition observations, as seen in Figures 7(a), 8(a) and 9. We base our suggested $\tilde{\theta}_{j}$’s on cropping the competition observations, as seen, e.g., in Figure 7(c), but this is far from the only approach. Shrinking, which we also investigated, might give a more realistic set of $\tilde{\theta}_{j}$’s, but it also requires an extra parameter. For the examples presented here shrinking gave similar results to cropping, and are therefore not included.
An alternative to the expected value of the observed maximum is the upper limit of the 95% CI. The latter is more in line with common interpretation of the uncertainty of observed results for research, but which to choose will depend on the aim of the analysis. We demonstrate the impact in the Obesity Risk example, reported in Table 3.
Not every $\hat{\theta}^{\,max}$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
@@ 5.1.1 Multi-Class Prediction of Obesity Risk
Multi-Class Prediction of Obesity Risk is a competition in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (10.34740/kaggle/dsv/7009925). The test set size was $n=13,840$, and the number of participating teams were $m=3,498$ (excluding teams with accuracy lower than chance). The best performing team had $\hat{\theta}^{\,\max}=0.9116$. Figure 7(a) shows the competition observations, $\hat{\theta}_{j}$’s, zoomed in at the $3,081$ best performing teams. The 95% (not multiplicity-corrected) CI for $\hat{\theta}^{\,max}$ is $(0.9067,0.9162)$ and is shown as the red line in Figure 7(a), the lower bound enveloping $920$ of the competition observations. This should raise suspicion that the right tail seen in Figure 7(a) can be the product of multiplicity.
(a) The competition observations
(b) A simulated realisation
(c) Cropped competition observations
(d) A simulated realisation
@@
To demonstrate the impact of ignoring multiplicity, let $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ be equal to the competition observations and calculate the expectation of the maximum value of a realisation. We use the set-up from Section 4.3 with correlation coefficient $\rho=0.6$, and random $Y_{0}$ with $\theta_{0}=\hat{\theta}^{\,max}$. Note that competition observations with low values are excluded in accordance with Eq. 12, but they are not influencing the maximum.
A realisation is shown in Figure 7(b). The upper tail, whose tip represents the maximum observed value, is stretched according to the variances of the $Y_{j}$s.
The estimated expected maximum is $E(\hat{\tilde{\theta}}^{\,max})=0.9159$, far above the observed maximum. The 95% CI is calculated by $B=1,000$ bootstrap samples of size $m$ drawn from the competition observations, and then the simulation as described above is conducted for each bootstrap sample. The values are reported in Table 3. The lower bound exceeds $\hat{\theta}^{\,\max}$, and we can conclude that the maximum observed value is an unlikely candidate for state-of-the-art.
Perhaps a more interesting problem is finding a set $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that can result in realisations similar to the competition observations in terms of maximum observed value, and let the maximum $\tilde{\theta}_{j}$ be the state-of-the-art candidate.
To obtain a set $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ where the expected maximum is the same as the maximum of the competition observations, $E(\hat{\tilde{\theta}}^{\,\max})=\hat{\theta}^{\,\max}$, we crop the competition observations, as seen in Figure 7(c), and where $E(\hat{\tilde{\theta}}^{\,\max})$ is calculated from simulations as described above. Figure 7(d) shows one realisation of the cropped distribution, where the maximum observed value is close to $\hat{\theta}^{\,max}$. By cropping at $0.9062$ we obtain $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$, and this serves as a new candidate for the state-of-the-art.
It can be argued that the new SOTA candidate is at risk of being a optimistic estimate, since the competition observations are from only one test set. The upper bound of the 95% CI can mitigate such a risk, and the competition observations are cropped to fit that criterion. The CI is calculated by bootstrapping, as described above. The results are shown in Table 3.
@@
Approach $\hat{\theta}_{SOTA}$ 95% CI $E(\max_{j}\theta_{j})$ teams
no mult. $\hat{\theta}^{\,\max}$ (0.9067,0.9162) - -
Kaggle accuracies $\hat{\theta}^{\,\max}$ (0.9128,0.9191) 0.9159 -
cropping 0.9062 (0.9084,0.9148) $\hat{\theta}^{\,\max}$ 1134
0.9026 (0.9050,$\hat{\theta}^{\,\max}$) 0.9083 1844
Table 3: The 95% CI and the expectation and for the SOTA estimator for different estimation strategies. The bootstrap standard deviations are in the order of magnitude $10^{\text{-}5}$-$10^{\text{-}4}$, and are therefore not displayed.
Table 3 demonstrates that taking the effect of multiplicity into account changes how results are perceived dramatically. A more likely SOTA value is $0.9062$, and even though its reduction from best performance is only about $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around one third of the teams have observed performance above the new SOTA. The more conservative approach using the upper bound of the 95% CI to estimate SOTA gives an even more dramatic result, and either way the impact is substantial.

### v5

@@ 4.1 Non-identical, independent classifiers
We are interested in generalising Eq. 3.3 to the non-identical case, and we start by noting that we now have $P_{xj}=P(X\leq x|\theta_{j})$, and we seek the sum of those so that we can find the proper expression for
$F(z|\mathbf{\Theta})=1-P(C_{z}=0|m,n,\mathbf{\Theta}).$
The Poisson Binomial distribution, see, e.g., Wang1993, describes the sum of $m$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{m}$. The pmf for $K$ number of successes is
@@
where $G_{k}$ is all subsets of size $k$ that can be selected from {1, 2, …, m}.
We then get the cdf for non-identical classifiers
$F(z|\mathbf{\Theta})$ $=1-\sum\limits_{{A\in G_{0}}}\prod\limits_{{i\in A}}P_{zi}\prod\limits_{{j\in A^{c}}}(1-P_{zj})=1-\prod\limits_{j=1}^{m}\big{(}1-P_{zj}\big{)}.$ (4)
@@
$f(z|\mathbf{\Theta})$ $=\prod\limits_{j=1}^{m}\big{(}1-P_{z\text{-}1j}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-P_{zj}\big{)}$ (5)
$=\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z\text{-1}}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}.$
@@
We use the same parameters as in Section 3.4, replacing $\theta=0.90$ with the vector $\mathbf{\Theta}$, where $\min(\mathbf{\Theta})=0.875,\,\max(\mathbf{\Theta})=0.90$ and equally spaced elements. The cdf and pmf are displayed in Fig. 5(a) and Fig. 5(b).
Cdf and pmf for non-identical classifiers.
(a) Cdf $F(z)$
(b) Pmf $f(z)$
Figure 5: The distribution and the probability function of the sample maximum estimator for non-identical $\theta$s, where $\min(\mathbf{\Theta})=0.875,\,\max(\mathbf{\Theta})=0.90$ and equally spaced elements.
The upper bound of the 95% CI is $0.9177$. The $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ is $0.9129$, with standard deviation of $0.0021$. Comparing this to the results in Section 3.4, we see that the bias is smaller and the variance larger. This is because the classifiers at the lower end does not contribute to the sample maximum, and the reduced number of contributing classifiers gives a higher variance. The exact reduction/increase in bias and variance depends on $\mathbf{\Theta}$, and can at its most extreme cases completely eradicate the multiplicity effect, as we shall see in Section 5.1.2. Figure 4(a) shows how $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ changes with $\min(\mathbf{\Theta})$ for equally spaced elements with $\max(\mathbf{\Theta})=0.90$.
What we see here is that with non-identical classifiers the bias decreases rapidly at first as $\min(\mathbf{\Theta})$ moves away from $\max(\mathbf{\Theta})$, but the rate at which this happens also decreases rapidly. This means that taking into account that classifier performances are heterogeneous is important, and small variations from identicality can have great impact on SOTA estimation. In Section 5, we demonstrate for other $\mathbf{\Theta}$s, based on observations.
4.2 Identical, dependent classifiers
Even though the different teams work independently, their resulting classifiers are not necessarily statistically independent. A classifier’s ability to perform better than random guesses is due to the correlations between features and labels in the training set. In their training phase, different classifiers will pick up on many of the same features and correlations, regardless of which classifier they train. In addition, many teams will use different versions of related models, often pre-trained on the same data set. This scenario can be modelled by conditional independency.
Let $Y_{0}$ be a Bernoulli variable with $P(Y_{0}=1)=\theta_{0}$. Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly with $P(Y_{j}=1)=\theta_{j}$, and let $j=1,\ldots,m$ be the $m$ classifiers contributing to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$. Boland et al. (1989) described dependency in the majority systems context, where $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$. The observations from $Y_{0}$ does not contribute directly to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, but can be seen as, e.g., the influence that a common training set has on the classifiers.
The binomial distribution for dependent Bernoulli trials is not an easy problem, see, e.g., Ladd (1975) and Hisakado et al. (2006). The Bernoulli trials can be simulated, see Boland et al. (1989) for conditional distributions. We give an example below for the case where $\rho_{0j}=\rho_{0}$ and $\theta_{j}=\theta$ for all classifiers. We set $\rho_{0}$ based on the work of Mania et al. (2019).
Mania et al. (2019) calculated model similarity, defined as the probability of two classifiers giving the same output, which is equivalent to
$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}$ (6)
for dependent models with correlation $\rho_{jk}$. The average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$, and the calculated similarity of $0.85$ gives $\rho=0.36$ if we assume the same correlation for all classifiers. We have that $\rho_{jk}=\rho_{0j}^{2}$, and hence our proposed correlation $\rho_{0}$ is set to $0.6$. Note that, for $\theta=0.90$, the similarity as described by Mania et al. (2019) is $0.82$ for two independent classifiers. When the probability of correct prediction is high, there is little room for variation between classifier outcomes.
Let $P(Y_{j}=1)=\theta=0.90$ for all classifiers, and $P(Y_{0}=1)=\theta=0.90$. From our simulations, we get $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9140$, a decrease in bias compared to the independent and identical classifiers. However, the variance has increased due to the random nature of $Y_{0}$, so the upper 95% CI is essentially the same. See, Table 2 for exact numbers.
To give an idea of how the randomness of $Y_{0}$ influences the variance, we fix the $y_{0}$’s with exactly $(n-x)/n=0.90$ correct predictions. The expected value is the same as with random $Y_{0}$s, but the standard deviation has now decreased from $0.0035$ to $0.0015$. In the real world, the dependencies have both random and fixed sources. A random source would, e.g., be the training set and the test set, random samples from an underlying distribution. A fixed source can be, e.g., fixed training conditions, such as pre-trained models.
For $\rho_{0j}=0$, we have independent classifiers, the same scenario as in Section 3.4, and for $\rho_{0j}=1$, we have $m$ classifiers producing the same output, and multiplicity is not a concern. For the whole range, see the red curve in Figure 4(b).
What we see here is that the correlation coefficient value in a model has a great impact on the estimated bias if the correlation is substantial. This can explain that in mature fields, where most competing teams will use variants of the same model, the multiplicity will not impact the overestimation of SOTA as much as in young fields where the competing teams have more diverse models. This aligns with what is often observed in science: overoptimistic results in the beginning, and more sober as the field evolves.
@@ 4.3 Non-identical, dependent classifiers
The most complex, but also the most realistic case, is where classifiers are non-identical and dependent. An analytical solution is not known Hisakado et al. (2006), but simulations are straightforward. The results of Boland et al. (1989) can be extended, and we have that
@@
Note that, since $P\leq 1$, it follows from Eq. 8 that
@@
and similarly for $0\leq P$ and Eq. 7.
We get an upper bound of the $95\%$ CI of $0.91733$ and an expected value of $0.9101$, which are lower than both the dependent and the non-identical model separately. The standard deviation is $0.0036$, which is higher than the two other models. This sums of up much of Table 2, where more complex, and therefore more realistic, models reduces the bias but increases the variances.
Figure 4(b) shows the impact of correlation for non-identical classifiers. What we see is that the bias decreases slowly for low correlations, but that the rate increases as $\rho_{0}$ increases. Importantly, Eq. 9 restricts the discrepancy between $\min(\mathbf{\Theta})$ and $\max(\mathbf{\Theta})$, and the graphs are right-truncated. From Fig. 3(a) we observe that the impact from number of classifiers flattens off, instead of restricting $\rho_{0}$ in our model, we can exclude classifiers by performance without large effect on the estimated bias. That is, restricting $\min(\mathbf{\Theta})$ instead of $\rho_{0}$. What we also observe in Fig. 4(a) is that change in $\min(\mathbf{\Theta})$ has less impact on the lower end of the scale, so this also supports a decision of excluding teams with low performance without impacting the bias-estimation too much.
@@ 5 Real world examples
As we have shown above with simulated data, the sample maximum is an overoptimistic estimate for SOTA. These results are based on known SOTA, but faced with real world data the problem to be solved is of opposite nature: The sample $\max(\hat{\Theta})$ is known, but the true SOTA must be estimated. In the public competition or public data set context there are many factors that come in to play, making the bias calculations everything but straightforward.
@@
In a public competition with $m$ participating teams, let $\hat{\Theta}=\{\hat{\theta}_{1},\hat{\theta}_{2},\ldots,\hat{\theta}_{m}\}$, be the estimated probability of success for each classifier based on the performance on an independent test sample of size $n$. The the winning team is identified by the sample maximum, $\max(\hat{\Theta})$. Among the different scenarios presented in Section 4, the non-identical dependent is the most realistic model. Whereas in Table 2 we report $E(\hat{\theta}^{\,max}_{SOTA})$ from known $\Theta$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\Theta}$. A straightforward approach is to find $\Theta^{\prime}=\{\theta^{\prime}_{1},\theta^{\prime}_{2},\ldots,\theta^{\prime}_{m}\}$ that gives $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$. That is, find a set of $\theta$’s so that the expected value of its sample maximum on a test set of size $n$ is equal to the competition sample maximum. The SOTA estimate is then $\max(\Theta^{\prime})$. We use a set-up where the simulated $\hat{\theta}^{\prime}_{j}$s are the mean values of $n$ realisations of $Y_{j}$, where $P(Y_{j})={\theta}^{\prime}_{j}$. The simulation is repeated $10,000$ times, and for each repetition the observed maximum value is recorded. $E(\max(\hat{\Theta}^{\prime}))$ is the mean value over the repetitions.
In the studies in Section 4, the $\theta_{1},\ldots,\theta_{m}$ were chosen to be equally spaced, but this does not align with the real world sample estimates, as seen in Figures 6(a), 7(a) and 8. We base our suggested ${\theta}^{\prime}_{j}$’s on cropping the sample estimates, as seen, e.g., in Figure 6(c), but this is far from the only approach. Shrinking, which we also investigated, might give a more realistic set of ${\theta}^{\prime}_{j}$’s, but it also requires an extra parameter. For the examples presented here shrinking gave similar results to cropping, and are therefore not included.
An alternative to the expected value of the sample maximum is the upper limit of the 95% CI. The latter is more in line with common interpretation of the uncertainty of observed results in research, but which to choose will depend on the aim of the analysis. We demonstrate the impact in the Obesity Risk example, reported in Table 3.
Not every $\max(\hat{\Theta})$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
@@ 5.1.1 Multi-Class Prediction of Obesity Risk
Multi-Class Prediction of Obesity Risk is a competition in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (Palechor and de la Hoz Manotas, 2023). The test sample size was $n=13,840$, and the number of participating teams were $m=3,498$ (excluding teams with accuracy lower than chance). The maximum sample estimate was $\max(\hat{\Theta})=0.9116$. Figure 6(a) shows the sample estimates, $\hat{\theta}_{j}$’s, zoomed in at the $3,081$ classifiers with sample estimates above $0.88$. The 95% CI for $\hat{\theta}=0.9116$ is $(0.9067,0.9162)$ and is shown as the red line in Figure 6(a), the lower bound enveloping $920$ of the sample estimates. This should raise suspicion that the right tail seen in Figure 6(a) can be the product of multiplicity.
(a) Sample estimates
(b) A simulated realisation from a)
(c) Cropped sample estimates
(d) A simulated realisation from c)
@@
To demonstrate the impact of ignoring multiplicity, let ${\theta}^{\prime}_{1},\ldots,{\theta}^{\prime}_{m}$ be equal to the sample estimates and calculate the expected value of the sample maximum, $E(\max(\hat{\Theta}^{\prime}))$. We use the set-up from Section 4.3 with correlation coefficient $\rho_{0}=0.6$, and $P(Y_{0})=\theta_{0}=\max({\Theta}^{\prime})$. Note that sample estimates with low values are excluded in accordance with Eq. 9, but their influence on the expected maximum is negligible.
A realisation is shown in Figure 6(b). The upper tail, whose tip represents the maximum sample estimate, is stretched according to the variances of the $Y_{j}$s. The estimated expected sample maximum is $E(\max(\hat{{\Theta}}^{\prime}))=0.9159$, far above the sample maximum in the competition. The 95% CI is reported in Table 3. The lower bound exceeds $\max(\hat{\Theta})$, and we can conclude that the sample maximum is an unlikely candidate for SOTA. The CI is calculated by $B=1,000$ bootstrap samples of size $m$ drawn from the sample estimates, and then the simulation as described above is conducted for each bootstrap sample.
Instead, we can find a set $\Theta^{\prime}$ that results in realisations, $\hat{\Theta}^{\prime}$, similar to the competition sample estimates in terms of sample maximum, and let $\max(\Theta^{\prime})$ be the SOTA candidate. To obtain a set $\Theta^{\prime}$ where $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$, we crop the competition sample estimates, as seen in Figure 6(c), and where $E(\max(\hat{\Theta}^{\prime}))$ is calculated from simulations as described above. Figure 6(d) shows one realisation of the cropped distribution, where $\max(\hat{\Theta}^{\prime})$ is close to $\max(\hat{\Theta})$. By cropping at $0.9062$ we obtain $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$, and this serves as a new candidate for the SOTA.
It can be argued that the new SOTA candidate is at risk of being a optimistic estimate, since the competition observations are from only one test sample. The upper bound of the 95% CI can mitigate such a risk, and the competition sample estimates are cropped to fit that criterion. The CI is calculated by bootstrapping, as described above. The results are shown in Table 3.
@@
Approach $\max(\Theta^{\prime})$ 95% CI $E(\max(\hat{\Theta}^{\prime}))$ teams
no mult. $\max(\hat{\Theta})$ (0.9067, 0.9162) - -
$\Theta^{\prime}=\hat{\Theta}$ $\max(\hat{\Theta})$ (0.9128, 0.9191) 0.9159 -
cropping 0.9062 (0.9084, 0.9148) $\max(\hat{\Theta})$ 1134
0.9026 (0.9050, $\max(\hat{\Theta})$) 0.9083 1844
Table 3: The 95% CI and the expected value for sample maximum, $\max(\hat{\Theta}^{\prime})$, for different SOTA estimation strategies, where SOTA is defined as $\max(\Theta^{\prime})$. The competition sample maximum is $\max(\hat{\Theta})=0.9116$. The bootstrap standard deviations are in the order of magnitude $10^{\text{-}5}$-$10^{\text{-}4}$, and are therefore not displayed. The rightmost column shows number of classifiers with sample performance above the estimated SOTA.
Table 3 demonstrates that taking the effect of multiplicity into account changes how results are perceived dramatically. The new SOTA value is $0.9062$, and even though its reduction from best performance is only about $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around one third of the classifiers have sample performance above the new SOTA. The more conservative approach using the upper bound of the 95% CI to estimate SOTA gives an even more dramatic result, and either way the impact is substantial.